# Bridging Operator Learning and Conditioned Neural Fields: A Unifying Perspective

### v1

@@ 4 Continuous Vision Transformer (CViT)
Finally, we use a small MLP network to project the output of the transformer backbone to the desired dimension.
Conditioning function of CViT.
The conditioning function takes the input function $u$ as an input and returns the latent conditioning variable $z_{L}$. We first construct a gridded representation of the input function $u$ yielding a 2+1 dimensional spatio-temporal data tensor $\mathbf{u}\in{\mathbb{R}}^{T\times H\times W\times D}$ with $D$ channels. We patchify our inputs into 3D tokens $\mathbf{u}_{p}\in{\mathbb{R}}^{T\times\frac{H}{P}\times\frac{W}{P}\times C}$ by tokenizing each 2D spatial frame independently using the same patch embedding process as in standard ViTs [49]. We then add (with broadcasting) trainable 1D temporal and 2D spatial positional embeddings to each token using broadcasting, as shown below:
Next, we use a temporal aggregation layer to reduce the token sequence length, and thereby the computational cost of subsequent self-attention layers [50]. This aggregation is performed using the Perceiver architecture put forth by Jaegle et al. [51]. Specifically, we learn a predefined number of latent input queries $\mathbf{z}\in{\mathbb{R}}^{1\times C}$ by applying a cross-attention module that maps the flattened visual features $\mathbf{u}_{f}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ to a latent tensor $\mathbf{z}_{agg}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times 1\times C}$ as
$\mathbf{z}_{agg}$ $=\mathbf{z}^{\prime}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}^{\prime}\right)\right),$
where $\mathbf{z}$ is trainable and initialized by a unit Gaussian distribution, and $\hat{\mathbf{z}}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times 1\times C}$ is obtained by tiling the latent query $\mathbf{z}$ $\left(\frac{H}{P}\times\frac{W}{P}\right)$ times. We then process the aggregated tokens $\mathbf{z}_{agg}$ using a sequence of $L$ pre-norm Transformer blocks [50, 52],
$\mathbf{z}_{0}$ $=\mathbf{z}_{agg},$
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L$
Figure 3: Continuous Vision Transformer (CViT): CViT builds a base field that consists of a trainable grid-based positional encoding of query coordinates $y$, followed by a standard transformer decoder [50]. The keys and values vectors of the base field are modulated via a global conditioning on the input function $u$, which is encoded by a vision transformer backbone.
Comparison to other transformer-based methods.
Here it is important to emphasize that the CViT base field performs cross-attention between each individual query point and the output tokens of the conditioning function. This is a subtle but significant difference from prior work, such as Oformer [53] and GNOT [54], where model predictions from different query locations are correlated because cross-attentions are computed for all queries simultaneously. In contrast, CViT vectorizes the cross-attention module across query coordinates via vmap in JAX [55]. Consequently, the model predictions corresponding to different query coordinates are independent of each other, thereby allowing us to build a well-defined conditioned neural field.
@@ 5 Experiments
We construct CViT models with different configurations, as summarized in Appendix Table 2. For all experiments, unless otherwise stated, we use a patch size of $8\times 8$ for tokenizing inputs. We also employ a base field with a single cross-attention Transformer block for all configurations. While we tested base fields with multiple layers, they did not yield improvements. The grid resolution is set to the spatial resolution of each dataset. The latent dimension of grid features is set to 512, and if it does not match the transformer’s embedding dimension, we align it using a dense layer. Besides, we use $\epsilon=10^{5}$ to ensure sufficient locality of the interpolated features. These choices are validated by extensive ablation studies, as illustrated in Figure 6. Full details of the training and evaluation procedures are provided in Appendix C.1.
@@
For all baselines, we report evaluation results from the original papers when applicable. When not applicable, we train and evaluate each model following the suggested settings in the respective paper. Details on the implementation of baseline models are provided in Appendix C.2.
@@ 5.1 Main Results
Here we provide our main results across three challenged benchmarks. The full details on the underlying equations, dataset generation and problem setup for each case are provided in the Appendix; see Section C.3, C.4 and C.5, respectively.
@@
The first benchmark involves predicting the transport of discontinuous waveforms governed by a linear advection equation with periodic boundary conditions. We make use of the datasets and problem setup established by Hoop et al. [61]. This benchmark evaluates our model’s capability in handling discontinuous solutions and shocks in comparison to popular neural operator models.
@@
The second benchmark involves a fluid flow system governed by the 2D shallow-water equations. This describes the evolution of a 2D incompressible flow on the surface of the sphere and it is commonly used as a benchmark in climate modeling. Here we adhere to the dataset and problem setup established in PDEArena [57], allowing us to perform fair comparisons with several state-of-the-art deep learning surrogates.
Table 5 presents the results of CViT against several competitive and heavily optimized baselines. Our proposed method achieves the lowest relative $L^{2}$ error. It is also worth noting that the reported performance gains are achieved using CViT configurations with a significantly smaller parameter count compared to the other baselines. This highlights the merits of our neural fields approach that enables the design of parameter-efficient models whose outputs can also be continuously queried at any resolution. Additional visualizations of our models are shown in Figure 9 and 10.
In addition, here we perform extensive ablation studies on the hyper-parameters of CViT. The results are summarized in Figure 6. We observe that a smaller patch size typically leads to better accuracy, but at higher computational cost. Moreover, model performance is significantly influenced by the type of coordinate embeddings used in the CViT base field. In order to ablate this choice, we consider two additional approaches to constructing these embeddings: (a) a small MLP, and (b) random Fourier features [62]. Our experiments reveal that the proposed grid-based embedding achieves the best accuracy, outperforming other methods by up to an order of magnitude. We also investigate the impact of the resolution of the associated grid. Our findings suggest that best results are obtained when the grid resolution matches the highest resolution at which the model is evaluated. Additionally, the CViT model shows sensitivity to the hyper-parameter $\epsilon$ used for computing interpolated features; too small $\epsilon$ values can degrade predictive accuracy. Finally, while we observe minor improvements by increasing the latent dimension of grid features, or the number of base field attention heads, the overall model performance remains robust to variations in these hyper-parameters; see Appendix Figure 8.
@@
[…]
Figure 5: Shallow-water equations benchmark: Representative CViT prediction of the vorticity field at 48hr lead time, and point-wise error versus the ground truth.
Figure 6: Ablation studies for CViT on the shallow-water equations benchmark. Convergence of test errors for: (a) different patch sizes; (b) different coordinate embeddings; (c) different resolutions of latent grids; (d) sensitivity on interpolating the CViT latent grid features, as controlled by the $\epsilon$ parameter. Results obtained using CViT-L with $16\times 16$ patch-size, varying each hyper-parameter of interest while keeping others fixed.
@@
Our last benchmark corresponds to predicting the evolution of a 2D buoyancy-driven flow, depicting a smoke volume rising in a bounded square domain. The underlying governing equations are the incompressible Navier-Stokes equations coupled with a passive scalar transport equation. For a fair comparison with various neural operators and other state-of-the-art deep learning models, we use the dataset generated by PDEArena [57], and follow the problem setup reported in Hao et al. [60].
@@
Our main results are summarized in Table 7, indicating that CViT outperforms all baselines we have considered. Notably, the previous best result (DPOT-L (Fine-tuned)) is achieved with a much larger transformer-based architecture, which involves pretraining on massive diverse PDE datasets followed by fine-tuning on the target dataset for 500 epochs. Furthermore, CViT exhibits remarkable parameter efficiency. CViT-small, with only 13M parameters, achieves accuracy comparable to the largest pretrained model (1.03B). Additionally, CViT shows good scalability; as the number of parameters increases, performance improves. These findings strongly support the effectiveness our approach in learning evolution operators for complex physical systems. Additional results and sample visualizations are provided in Figure 7, as well as Figure 11, 12 and 13 in the Appendix.
[…]
tablePerformance on the Navier-Stokes benchmark. All mode (128 × 128 resolution for both training and testing). Model # Params Rel. $L^{2}$ error ($\downarrow$) FNO 0.5 M 9.12 % FFNO 1.3 M 8.39 % GK-T 1.6 M 9.52% GNOT 1.8 M 17.20 % Oformer 1.9 M 13.50 % DPOT-Ti 7 M 12.50 % DPOT-S 30 M 9.91 % DPOT-L (Pre-trained) 500 M 7.98 % DPOT-L (Fine-tuned) 500 M 2.78 % DPOT-H (Pre-trained) 1.03 B 3.79 % CViT-S 13 M 3.75 % CViT-B 30 M 3.18 % CViT-L 92 M 2.35 %
@@
Figure 7: Navier-Stokes benchmark. Representative CViT predictions of the passive scalar field, and point-wise error against the ground truth.
@@ 6 Discussion
This work uncovers a novel connection between operator learning architectures and conditioned neural fields from computer vision. By reformulating popular operator learning models as neural fields with specific conditioning mechanisms, we provide a unifying perspective for understanding and comparing these architectures. Motivated by these findings, we introduce the Continuous Vision Transformer (CViT), a new operator learning architecture that leverages the power of vision transformers and cross-attention to modulate a transformer base field, achieving state-of-the-art accuracy on challenging benchmarks in climate modeling and fluid dynamics. These contributions demonstrate the potential of adapting advanced computer vision techniques to develop more flexible and accurate machine learning models for physical sciences, bridging the gap between the two fields and paving the way for further research in this direction. The broader impact of this work lies in its potential to fuel the development of resolution-invariant foundation models for accelerating scientific discovery and enable more efficient and accurate simulations of complex physical systems, with applications ranging from climate modeling to engineering design.
@@
Despite its promising performance, the proposed CViT architecture faces several limitations and open challenges . Firstly, like most transformer-based models, CViT’s self-attention blocks scale quadratically with the number of tokens, posing computational challenges for high-dimensional or high-resolution data. Recent advancements in computer vision [63] are tackling these scalability issues, and these ideas can be transferred to our setting. Secondly, transformer-based models like CViT require large training datasets for optimal performance, and may struggle in data-scarce scenarios without careful regularization. Leveraging CViT’s ability to reconstruct continuous output functions, physics-informed loss functions [64, 65] can be used for data-efficient training and regularization. Lastly, while the experiments presented here treat input functions as a sequence of images, many practical applications involve complex geometric structures and diverse input modalities like meshes and point clouds. Addressing these cases requires developing appropriate tokenization schemes, and recent work in [66, 67] has taken steps in this direction.
Potential Negative Societal Impacts.
Our contributions enables advances in accelerating the modeling and emulation of physical systems. As with any tool that furthers our understanding and ability to predict the outcomes of complex systems, there may be ill-intentioned use cases, but we do not expect any specific negative impact from this work.
Acknowledgement
We would like to acknowledge support from the US Department of Energy under the Advanced Scientific Computing Research program (grant DE-SC0024563). We also thank the developers of the software that enabled our research, including JAX [55], Matplotlib [68], and NumPy [69].

### v2

@@ 3.1 Architecture Description
Finally, a small MLP network projects the output to the desired output dimension.
Theoretical Insights.
The theoretical analysis of Lipschitz constants for different coordinate embeddings (as detailed in Appendix E) provides crucial insights that inform the design choices in CViT and have broader implications for neural operator architectures. Our findings reveal that the choice of coordinate embedding significantly influences the decoder’s capacity to learn complex, high-frequency functions.
For conventional MLP decoders with linear coordinate embeddings, the Lipschitz constant is approximately 1 at initialization, which does not enhance the network’s ability to capture fine-scale features, thereby leading to prediction susceptible to spectral bias (Rahaman et al., 2019). In contrast, random Fourier features can increase the Lipschitz constant through two mechanisms: by increasing the standard deviation of the sampling distribution or by increasing the network width. This observation not only explains the effectiveness of Fourier features in addressing spectral bias, but also highlights a potential trade-off between network size and the frequency range that can be captured.
Our proposed grid-based coordinate embedding in CViT offers a more direct and controllable approach to increasing the network’s Lipschitz constant. By adjusting the interpolation parameter $\beta$, we can tune the model’s ability to capture high-frequency components in the target functions. This provides a clear advantage as it allows for more precise control over the model’s spectral properties without necessarily increasing model complexity.
CViT vs other Transformer-based Approaches.
To further highlight the key innovations of CViT, we compare it with other popular transformer-based operator learning methods:
OFormer (Li et al., 2022): OFormer uses an encoder-decoder structure. Its encoder combines a point-wise MLP and stacked self-attention blocks to process input functions, while employing random Fourier Features (Tancik et al., 2020) for query coordinate encoding. The model generates latent embeddings through cross-attention at sampled query locations. In the decoder, OFormer uses recurrent MLPs for dynamic propagation instead of traditional masked attention. Unlike CViT, it doesn’t use ViT’s patch embedding to tokenize input functions, potentially limiting its ability to capture complex spatial dependencies.
GNOT (Hao et al., 2023): GNOT is a transformer-based framework for learning operators that addresses the challenges of multiple input functions and irregular meshes. It designs a heterogeneous normalized attention layer, providing an encoding interface for various input functions and additional prior information. Additionally, GNOT features a geometric gating mechanism, described as a soft domain decomposition approach to multi-scale problems. Unlike CViT, GNOT uses an plain MLP to process query coordinates, which may potentially suffer from spectral bias (Rahaman et al., 2019).
DPOT (Hao et al., 2024): DPOT is a transformer-based architecture designed for pretraining on multiple PDEs. It introduces a new auto-regressive denoising strategy and a novel model architecture based on Fourier attention. Unlike OFormer, GNOT, and CViT, DPOT does not explicitly take query coordinates as inputs, which implies that it cannot be evaluated at arbitrary locations.
MPP (McCabe et al., 2023): MPP is a large-scale PDE surrogate model that directly employs a Vision Transformer as its backbone. Similar to DPOT, it does not take query coordinates as inputs, thereby losing flexibility in diverse output representations.
@@ 4 Experiments
We construct CViT models with different configurations, as summarized in Table 1. For all experiments, unless otherwise stated, we use a patch size of $8\times 8$ for tokenizing inputs. We also employ a decoder with a single cross-attention Transformer block for all configurations. While we tested decoders with multiple layers, they did not yield improvements. The grid resolution is set to the spatial resolution of each dataset. The latent dimension of grid features is set to 512, and if it does not match the transformer’s embedding dimension, we align it using a dense layer. Besides, we use $\beta=10^{5}$ to ensure sufficient locality of the interpolated features. These choices are validated by extensive ablation studies, as illustrated in Figure 4. Full details of the training and evaluation procedures are provided in Appendix F.1.
Table 1: Details of Continuous Vision Transformer model variants.
Model Encoder layers Embedding dim MLP width Heads # Params
CViT-S 5 384 384 6 13 M
CViT-B 10 512 512 8 30 M
CViT-L 15 768 1536 12 92 M
@@
For all baselines, we report evaluation results from the original papers when applicable. When not applicable, we train and evaluate each model following the suggested settings in the respective paper. Details on the implementation of baseline models are provided in Appendix F.2.
@@ 4.1 Main Results
Here we provide our main results across three challenged benchmarks. The full details on the underlying equations, dataset generation and problem setup for each case are provided in the Appendix; see Section F.3, F.4 and F.5, respectively.
@@
The first benchmark involves predicting the transport of discontinuous waveforms governed by a linear advection equation with periodic boundary conditions. We make use of the datasets and problem setup established by Hoop et al. (de Hoop et al., 2022). This benchmark evaluates our model’s capability in handling discontinuous solutions and shocks in comparison to popular neural operator models.
@@
The second benchmark involves a fluid flow system governed by the 2D shallow-water equations. This describes the evolution of a 2D incompressible flow on the surface of the sphere and it is commonly used as a benchmark in climate modeling. Here we adhere to the dataset and problem setup established in PDEArena (Gupta & Brandstetter, 2022), allowing us to perform fair comparisons with several state-of-the-art deep learning surrogates.
Table 2: Performance for 5-step roll-out predictions in the shallow-water equations benchmark.
Model # Params Rel. $L^{2}$ error ($\downarrow$)
DilResNet 4.2 M 13.20%
@@
FNO 268 M 3.97%
U-F2Net 344 M 1.89%
UNO 440 M 3.79%
CViT-S 13 M 4.47%
CViT-B 30 M 2.69%
CViT-L 92 M 1.56%
Table 2 presents the results of CViT against several competitive and heavily optimized baselines. Our proposed method achieves the lowest relative $L^{2}$ error. It is also worth noting that the reported performance gains are achieved using CViT configurations with a significantly smaller parameter count compared to the other baselines. This highlights the merits of our approach that enables the design of parameter-efficient models whose outputs can also be continuously queried at any resolution. Additional visualizations of our models are shown in Figure 3 and 10.
In addition, here we perform extensive ablation studies on the hyper-parameters of CViT. The results are summarized in Figure 4. As shown in Figure 4 (a), we observe that a smaller patch size typically leads to better accuracy, but at higher computational costs. Moreover, Figure 4 (b) demonstrates that the model performance is significantly influenced by the type of coordinate embeddings used in the CViT decoder. We compare our proposed grid-based embedding against a small MLP and random Fourier features (Tancik et al., 2020). The grid-based embedding achieves the best accuracy, outperforming other methods by up to an order of magnitude. This result is in agreement with the theoretical insights presented in Appendix E.
We also investigate the impact of the resolution of the associated grid. Figure 4 (c) reveals that the best results are obtained when the grid resolution matches the highest resolution at which the model is evaluated. As another direct empirical evidence of theoretical results in Appendix E, Figure 4(d) shows the CViT model’s sensitivity to the hyper-parameter $\beta$ used for computing interpolated features; too small $\beta$ values can degrade predictive accuracy. Finally, in Appendix Figure 9, we observe minor improvements by increasing the latent dimension of grid features, or the number of decoder attention heads. Hence the overall model performance remains robust to variations in these hyper-parameters.
Taken together, these findings not only justify our design choices for CViT but also provide general insights for the development of neural operators. They suggest that future research in this field should pay close attention to the interplay between spatio-temporal tokenization, coordinate representation, and the inherent multi-scale nature of physical systems. Such considerations could lead to even more effective architectures for a wide range of scientific machine learning tasks.
Figure 4: Ablation studies for CViT on the shallow-water equations benchmark. Convergence of test errors for: (a) different patch sizes; (b) different coordinate embeddings; (c) different resolutions of latent grids; (d) sensitivity on interpolating the CViT latent grid features, as controlled by the $\beta$ parameter. Results obtained using CViT-L with $16\times 16$ patch-size, varying each hyper-parameter of interest while keeping others fixed.
@@
Our third benchmark corresponds to predicting the evolution of a 2D buoyancy-driven flow, depicting a smoke volume rising in a bounded square domain. The underlying governing equations are the incompressible Navier-Stokes equations coupled with a passive scalar transport equation. For a fair comparison with various neural operators and other state-of-the-art deep learning models, we use the dataset generated by PDEArena (Gupta & Brandstetter, 2022), and follow the problem setup reported in Hao et al. (Hao et al., 2024).
@@
Our main results are summarized in Table 3, indicating that CViT outperforms all baselines we have considered. Notably, the previous best result (DPOT-L (Fine-tuned)) is achieved with a much larger transformer-based architecture, which involves pretraining on massive diverse PDE datasets followed by fine-tuning on the target dataset for 500 epochs. Here we also note that DPOT was pretrained and fine-tuned using a rollout loss, which directly optimizes for the multi-step prediction task on which the models are evaluated. In contrast, CViT was trained using a one-step loss without any rollout finetuning. This difference in training objective significantly favors DPOT for rollout evaluations across all benchmarks. Furthermore, CViT exhibits remarkable parameter efficiency. CViT-S, with only 13M parameters, achieves accuracy comparable to the largest pretrained model (1.03B). Additionally, CViT shows good scalability; as the number of parameters increases, performance improves. These findings strongly support the effectiveness our approach in learning evolution operators for complex physical systems. Additional results and sample visualizations are provided in Figure 5, as well as Figure 11 and 12 in Appendix.
Table 3: Relative $L^{2}$ error of rollout predictions on the Navier-Stokes and Diffusion-Reaction benchmark.
Model # Params Navier-Stokes Diffusion-Reaction
FNO 0.5 M 9.12 % 12.00 %
FFNO 1.3 M 8.39 % 5.71 %
GK-T 1.6 M 9.52% 3.59 %
GNOT 1.8 M 17.20 % 3.11 %
Oformer 1.9 M 13.50 % 1.92 %
DPOT-Ti 7 M 12.50 % 3.21 %
MPP-S 30M - 1.12 %
DPOT-S 30 M 9.91 % 3.79 %
MPP-L (Pre-trained) 400 M - 0.98 %
DPOT-L (Pre-trained) 500 M 7.98 % 2.32 %
DPOT-L (Fine-tuned) 500 M 2.78 % 0.73 %
DPOT-H (Pre-trained) 1.03 B 3.79 % 1.91 %
CViT-S 13 M 3.75 % 1.13 %
CViT-B 30 M 3.18 % 1.11 %
CViT-L 92 M 2.35 % 0.68 %
@@
Figure 5: Navier-Stokes benchmark. Representative CViT rollout predictions of the passive scalar field, and point-wise error against the ground truth.
Diffusion-Reaction equations.
Our final benchmark focuses on learning the density fields $\mathbf{u}=(u,v)$ governed by a coupled 2D diffusion-reaction equation (DR). This PDE, commonly used in chemistry, models substances reacting and diffusing across a spatial domain, illustrating chemical reaction processes and resulting in diverse spatial patterns. For a fair comparison, we use the dataset from PDEBench (Takamoto et al., 2022) and follow the problem setup described by Hao et al. (2024).
Table 3 presents our main results. Consistent with previous findings, CViT-L achieves the best performance across all baselines. CViT models outperform other approaches with comparable or fewer parameters, demonstrating good scalability and parameter efficiency. Representative rollout predictions are provided in Figure 6 and 13 (see Appendix). Our predictions show excellent agreement with the corresponding ground truth.
Interestingly, we note that the solution of Diffusion-Reaction equations appears simpler than that of the Navier-Stokes equation dataset, resulting in uniformly lower test errors for various baselines in the DR benchmark. In particular, OFormer outperforms GNOT in both examples, despite having comparable parameters. We postulate that this performance gain is due to the use of random Fourier features in OFormer, facilitating more efficient learning of complex functions. This observation further demonstrates the importance of coordinate embedding in designing neural operators and indirectly highlights the effectiveness of our proposed grid-based embedding.
Figure 6: Diffusion-reaction Benchmark. Representative CViT rollout prediction of the density field $u$, and point-wise error against the ground truth.
@@ 5 Discussion
This work introduces the Continuous Vision Transformer (CViT), a novel neural operator architecture that leverages advances in computer vision to address challenges in learning complex physical systems. CViT combines the strengths of vision transformers and continuous function representations, achieving state-of-the-art accuracy on challenging benchmarks in climate modeling, fluid dynamics and reaction-diffusion processes. Our approach demonstrates the potential of adapting advanced computer vision techniques to develop more flexible and accurate machine learning models for physical sciences. The key innovations of CViT include: (a) a vision transformer encoder for capturing multi-scale spatial dependencies in input functions, (b) a trainable grid-based positional encoding for flexible representation of query coordinates, and (c) a query-wise cross-attention mechanism ensuring consistent evaluation at arbitrary resolutions. These design choices ensure that CViT maintains the desirable properties of a well-defined conditioned neural field (as discussed in Appendix B), while offering improved flexibility and performance (as discussed in Appendix E).
Our empirical results across various PDE benchmarks demonstrate that CViT often outperforms larger models with fewer parameters and without extensive pretraining. This highlights the merits of our approach in designing parameter-efficient models that can be continuously queried at any resolution. The broader impact of this work lies in its potential to accelerate scientific discovery through more efficient and accurate simulations of complex physical systems, with applications ranging from climate modeling to engineering design.
@@
Despite its promising performance, the proposed CViT architecture can be further improved in multiple areas. Firstly, like most transformer-based models, CViT’s self-attention blocks scale quadratically with the number of tokens, posing computational challenges for high-dimensional or high-resolution data. Recent advancements in computer vision (Liu et al., 2021b) are tackling these scalability issues, and these ideas can be transferred to our setting to improve CViT’s efficiency for larger-scale problems. Secondly, transformer-based models like CViT generally require large training datasets for optimal performance and may struggle in data-scarce scenarios without careful regularization. To address this, future work could explore leveraging CViT’s ability to reconstruct continuous output functions in conjunction with physics-informed loss functions (Wang et al., 2021; 2023) for more data-efficient training and improved generalization. Lastly, while our current experiments treat input functions as sequences of images, many practical applications in scientific computing involve complex geometric structures and diverse input modalities such as unstructured meshes or point clouds. Addressing these cases requires developing appropriate tokenization schemes, and recent work in (Alkin et al., 2024; Pang et al., 2022) has taken steps in this direction. Extending CViT to handle such diverse input modalities represents an important avenue for future research.
Future Work.
Building on the success of CViT, several promising directions for future research emerge. One key area is the exploration of more efficient attention mechanisms to improve scalability to higher-dimensional problems. Additionally, investigating the potential of pre-training strategies, similar to those employed in DPOT (Hao et al., 2024), could further enhance CViT’s performance and generalization capabilities across diverse PDE systems. Finally, extending CViT to handle 3D problems and developing techniques for multi-resolution training and evaluation could significantly broaden its applicability in scientific computing domains.
Broader Impact.
The broader impact of this work lies in its potential to fuel the development of resolution-invariant foundation models for accelerating scientific discovery and enable more efficient and accurate simulations of complex physical systems, with applications ranging from climate modeling to engineering design.
Acknowledgements
We would like to acknowledge support from the US Department of Energy under the Advanced Scientific Computing Research program (grant DE-SC0024563). We also thank the developers of the software that enabled our research, including JAX Bradbury et al. (2018), Matplotlib Hunter (2007), and NumPy Harris et al. (2020).
Ethics Statement
Our contributions enable advances in accelerating the modeling and emulation of complex physical systems. This has the potential to significantly impact fields such as climate science, fluid dynamics, and materials engineering by providing more accurate and efficient simulation tools. While we do not anticipate specific negative impacts from this work, as with any powerful predictive tool, there is potential for misuse. We encourage the research community to consider the ethical implications and potential dual-use scenarios when applying these technologies in sensitive domains.
Reproducibility Statement
The code used to carry out the ablation experiments is provided as a supplementary material for this submission. If the paper is accepted, we plan on making our entire code-base publicly available on GitHub.